# Simulating Human-like Daily Activities with Desire-driven Autonomy

### v1

@@ 6 Experiment
In this section, we present a series of experiments conducted within a predefined environment to address three key research questions: (1) Does the Desire-driven Agent (D2A) generate more human-like daily activities compared to baseline approaches? (2) Can the activities generated by D2A satisfy its desires like humans? (3) How do the various components of our framework impact the performance and behaviors of D2A?
@@ 6.1 Environment
Our main experimental environment is the indoor environment, which consists of a house with four rooms—a kitchen, living area, bedroom, and bathroom—inhabited by a character named Alice. Alice has a detailed profile reflecting her priorities and characteristics across the 11 desire dimensions. The house is fully enclosed, and various items are distributed across the rooms to meet Alice’s needs, creating a comprehensive setting for testing agent behavior. The agent’s primary goal is to live comfortably and fulfill its desires by interacting with the available items. Additional details on the setup can be found in the Appendix A.1.
We also introduced a multi-agent social setting at a simulated outdoor environment, Central Park (See Appendix A.2). This environment hosts a lively “Enjoy Your Life” party featuring diverse facilities and activities. We added two background agents (See Appendix A.3 in this environment, providing agents with opportunities to address higher-level desires, especially desires related to social activities. This scenario enables us to validate the scalability of our framework in handling complex, dynamic, and socially rich environments.
We compare our Desire-Driven Agent with 3 baselines: ReAct (Yao et al. (2023)) agent reasons before taking actions; LLMob (Wang et al. (2024a)) generates activity plans by extracting motivations from the profile and producing the most likely activities accordingly; BabyAGI (Nakajima (2023)) maintains a prioritized task list and selects the next activity based on priority. To evaluate the effectiveness of the specific designs made for D2A, desire states $D^{t}$ are not provided to the baseline agents. Instead, they are given the goal, profile, or more descriptive habits (Appendix N) generated by the profile according to their respective methods. As a reference for human behavior, we also include a human-controlled agent, where the player is given the same desired value provided to the D2A. For all the four agent-based methods, we used LLaMA3.1-70B as the default backbone model for both the agents and the environment controller. We also conducted experiments on a stronger backbone model Qwen2.5-72B (Team, 2024) to discuss how the capacity of large language models affects the performance of D2A.
@@ 6.2 Evaluation Metrics
We use two primary metrics to evaluate the performance of our agent compared to the baselines: Likelihood and Dissatisfaction. By employing these metrics, we can assess the human-likeness as well as the desire-satisfying ability of the proposed agents.
@@
Likelihood measures how closely an agent’s behavior resembles that of a human. Given the difficulty of directly measuring human-likeness, we define likelihood by employing an external model to assess the win rate between activity sequences generated by two agents, determining which sequence appears more human-like. We conduct this ”Player vs Player” comparison across the activity sequences produced by the baselines and the D2A agent. Specifically, after obtaining all action sequences $A_{p}$ for agent $p$, represented as $[A^{1}_{p},\ldots,A^{N}_{p}]$, we randomly select $seq_{i}$, $seq_{j}$ from $[A_{i}]$, $[A_{j}]$, respectively, for a given pair of agents $i$, $j$. These sequences are then compared using GPT-4o. This process is repeated 100 times for each agent pair $i$ and $j$ ($\forall i\neq j$), and the win rate for each agent is calculated. The results are visualized using a heatmap.
@@
Dissatisfaction measures the discrepancy between the agent’s current desire values and its expected values. Dissatisfaction $\delta^{t}$ of step $t$ is defined as the relative difference between current numerical desire values $v^{t}$ and expected desire values $v^{*}$ at the end of simulation step $t$. We calculate $\delta^{t}$ after the agent takes the $t$-th activity and updates its values. The formula used is:
@@ 6.3.1 Does D2A Generate More Human-like Daily Activities?
We conducted 15 trials for each agent under identical initialization settings across all experiments. After gathering each agent’s activity sequences, we found differences in output styles (See Appendix W) that could influence the evaluation by GPT-4o. To mitigate this, we preprocessed all sequences using Llama3.1 70B for style-level rewriting. This ensured more consistent activity sequences across agents, allowing for a fair comparison in the subsequent evaluation.
We used the method mentioned in Section (6.2) to obtain the relative likelihood map of approaches. When evaluating the likelihood of the activity sequences, GPT-4o is instructed to take the following criteria into account: Naturalness: How well the activity aligns with the person’s innate abilities, habits, and environment. Coherence: How logically and seamlessly different actions or steps within a sequence fit together to achieve a goal. Plausibility: How reasonable, likely, or believable a sequence of actions is, given the circumstances, context, and known behavior patterns. GPT-4o was asked to consider all three criteria together and output a final determination of which sequence was more human-like. See Appendix Y.1 for the prompts.
@@
The results presented in Figure 3 indicate that the D2A agent outperforms other baseline agents. We validated through human experiments that GPT-4o’s evaluations exhibit a high level of consistency with human judges. (See Appendix 1)This superior performance may be attributed to the fact that D2A is motivated by desire, aligning with the theory of human needs. Guided by desire, D2A effectively simulates human-like activity sequences, demonstrating a level of naturalness due to the recognizable characteristics of its profile. Furthermore, coherence is maintained as D2A considers its current desires based on the activity proposal component, continuously engaging in specific actions to achieve its goal of living comfortably until its desires are fulfilled. In contrast, ReAct, as the foundational agent, operates on a goal-reasoning framework. Its reliance on reasoning for action selection ensures reasonable interactions with environmental elements, thereby enhancing plausibility. However, the absence of a structured plan diminishes both the coherence and naturalness of its actions. BabyAGI, on the other hand, generates subsequent tasks based on its profile and motivational factors and maintains the priority list based on the observation. Although this approach facilitates coherence, it limits the diversity, as each activity is based on the outcomes of the previous one. Consequently, the agent’s tendency to engage repetitively in similar behaviors (enhancing its spiritual satisfaction based on the profile (See Appendix T), may result in diminished performance compared to D2A due to a lack of action diversity. LLMob leverages both planning and motivation derived from its profile to create a broad activity space. However, since it is characteristic-oriented, it keeps performing activities related to the habits derived from its profile. As a result, there may be no connection between activities, reducing overall coherence. The comparison of action diversity can be found at Appendix D.1.
@@ 6.3.2 Can the activities generated by D2A satisfy its desires like humans?
(a) The Averaged Dissatisfaction result over time steps of the Random-8-Steps Experiment.
(b) The Averaged Dissatisfaction result over time steps of the Fixed-12-Steps Experiment.
@@
To further evaluate and explain D2A’s activity generation capability, we conducted case studies in both indoor and outdoor environments. These studies focused on two aspects: (1) how actions are driven by Desire, and (2) the diversity of generated actions. Specifically, we examined how the Dissatisfaction between current and expected Desire status influences action selection (See Appendix C for further details). Our results also demonstrate that D2A’s desire-driven approach enables the exploration of a wider range of actions in environments with diverse action possibilities. More details and analysis can be found in Appendix D. These empirical results support our hypothesis and confirm that desire-driven modeling plays a crucial role in enabling D2A to make more informed decisions, ultimately producing activity sequences that closely mimic those of humans.
6.3.3 Ablation Study
@@
(d) Base Model
Figure 5: Ablation study regarding (a) Qualitative Value Description, (b) Searching Width of the Desire-Driven Planner. (c) Additional Multi-step Plan Component and (d) Different Base model. Please note that due to the significant time required to run 42 steps, neither of the two 42-step graphs includes a standard deviation.
@@
As described in Section 5.3, our D2A doesn’t contain a multi-step planning module that plans the next few steps’ activities ahead, like some other approaches in consideration of the dynamic nature of desires. We performed an ablation study by adding Concordia’s multi-step planning component, which updates the plan for the next few hours based on current desire states. As shown in Figure 5(c), the absence of the multi-step planning component allows D2A to exhibit a better ability to respond to immediate desire states. This may be because when the multi-step planning component is present, D2A tends to adhere to the pre-defined activities in the plan, which may limit its responsiveness to real-time changes in desire states. Furthermore, the plan is scheduled in hourly increments, and D2A cannot effectively predict which desires will experience the greatest shortfall in the next several hours. This limitation may contribute to the slight decline in performance observed when the plan component is included.
Finally, to evaluate the adaptability of our framework across different large language models, we tested it with Qwen 2.5 72B (Yang et al., 2024), which can also be integrated through the Ollama interface. As demonstrated in Figure 5(d), our framework, when running on the Qwen 2.5 72B base model, successfully provides desire-driven action guidance for the agent.
@@ 7 Conclusion and Discussion
We introduced the Desire-driven Autonomy (D2A) framework to simulate human-like daily activities by aligning agent behavior with intrinsic motivations. Drawing inspiration from Maslow’s theory of needs, our framework models agents’ desires across multiple dimensions and allows them to autonomously generate and evaluate activities to fulfill these desires. Through extensive simulated experiments, we demonstrated that D2A produces more natural, coherent, and plausible activity sequences compared to baseline methods, including ReAct, BabyAGI, and LLMob. Our experimental results support the hypothesis that modeling intrinsic desires enables agents to make more human-like decisions and behaviors. Moreover, the D2A framework offers promising avenues for further development, as it has shown adaptability with different large language models and potential for enhanced performance with stronger backbones and broader activity proposals. The results provide a solid foundation for future work in desire-driven AI systems for simulating complex, human-like interactions and behaviors.
@@
Acknowledgements
This work was supported by the National Science and Technology Major Project (2022ZD0114904), NSFC-6247070125, NSFC-62406010, the State Key Lab of General Artificial Intelligence at Peking University, Qualcomm University Research Grant, and Wuhan East Lake High-Tech Development Zone, National Comprehensive Experimental Base for Governance of Intelligent Society. We thank Wei Wang, Junqi Wang, Prof. Jiayu Zhan, and Prof. Song-Chun Zhu for their helpful discussion in our early work.
@@ Appendix B Consistency between Human Annotators and GPT-4o Evaluations
To validate the robustness of GPT-4o’s evaluations, we randomly selected 19 pairs of activity sequences for pairwise comparison (16 pairs from the indoor scenarios and 3 pairs from the outdoor scenarios) and created three questionnaires, each assigned to 8 human annotators. The human annotators were asked to judge which sequence in each pair appeared more human-like. Based on the level of agreement among the annotators, we categorized the 18 samples into three groups: over 75% agreement on one side—high consensus among humans; 50.1%–74.9%—moderate human preference; and exactly 50%—indicating humans found it difficult to distinguish between the two. We also provided the same pairs to GPT-4o, using our evaluation promptY.1 and both sequences as context, and asked it to determine which sequence appeared more human-like. The consistency rate between human evaluations and GPT-4o’s assessments is summarized as follows.
@@ C.1 Outdoor environment
At the beginning of the game, when she selected action for the next 20 minutes in Step 1, the Desire components output descriptive sentences about her current Desire status (e.g., ”Completely isolated, lacking any meaningful social connections”). These insights guided her to engage in social activities; for example, she joined a group discussion. Consequently, her Dissatisfaction with social connectivity, recognition, and sense of control decreased. Since we did not allow the agent to engage in detailed conversations, she did not have the opportunity to showcase herself, which kept her sense of superiority unchanged.
In Step 2, her sleepiness began to increase, and she still felt some Dissatisfaction regarding her sense of control. Therefore, she chose to take a short stroll, which could slightly alleviate her sleepiness. However, as a result of this short stroll, her thirst increased, prompting Alice to visit the refreshment station.
In Step 9, as her dissatisfaction with recognition and social connectivity grows, she decides to visit the party’s photo booth to gain some recognition and establish social connections.
Figure 6: The dissatisfaction changes at each step; the action $i$ will cause the dissatisfaction of step $i+1$ to change
Figure 7: The desire changes at each step; the action $i$ will cause a change in dissatisfaction in step $i+1$. We define some desires (Hunger, Thirst, Sleepiness) such that lower values are better. For better visualization, we adjust the magnitude so that higher values represent better outcomes.
@@ C.2 Indoor environment
Figure 8: The Dissatisfaction changes in each step, the action $i$ will caused the dissatisfaction change in step $i+1$
Figure 9: The desire changes at each step; the action $i$ will cause a change in dissatisfaction in step $i+1$. We define some desires (Hunger, Thirst, Sleepiness) such that lower values are better. For better visualization, we adjust the magnitude so that higher values represent better outcomes.
Initially, in Alice’s dissatisfaction, thirst, passion, and social connectivity are relatively significant. Therefore, she decides to first enhance her more basic desires (hunger and thirst, and regain some energy). So, she makes a hearty breakfast (Alice heads to the kitchen to prepare a nutritious breakfast, consisting of scrambled eggs, whole-grain toast, and a glass of freshly squeezed orange juice). This also successfully boosts her corresponding Desire component. Since we are using an action selection method similar to MCTS, the other two action choices are: take a refreshing shower in the bathroom (cleanliness, health, comfort) and meditate for a few minutes (reduce loneliness, increase sense of safety, and joyfulness). In the next step, she hopes to reduce her dissatisfaction with cleanliness. Therefore, she chooses to take a shower using the aforementioned method, but since she is not currently in the bathroom, she heads there first. As a result, her dissatisfaction with cleanliness does not decrease at this step. However, once she finishes showering in the bathroom, her cleanliness reaches satisfaction. After fulfilling most of her physical dimensions, she decides to address her Desire for the mental dimension (social connectivity). Consequently, she tries to call family or friends, which reduces her dissatisfaction with social connectivity.

### v2

@@ 6 Experiment
In this section, we present a series of experiments conducted within a predefined environment to address three key research questions: (1) Does the Desire-driven Agent (D2A) generate more human-like daily activities compared to baseline approaches? (2) Can the activities generated by D2A satisfy its desires like humans? (3) Can D2A perform well across different environments? (4) How do the various components of our framework impact the performance and behaviors of D2A?
@@ 6.1 Environment
Our main experimental environment is the indoor environment, which consists of a house with four rooms—a kitchen, living area, bedroom, and bathroom—inhabited by a character named Alice. Alice has a detailed profile reflecting her priorities and characteristics across the 11 desire dimensions. The house is fully enclosed, and various items are distributed across the rooms to meet Alice’s needs, creating a comprehensive setting for testing agent behavior. The agent’s primary goal is to live comfortably and fulfill its desires by interacting with the available items. Additional details on the setup can be found in the Appendix A.1. We also introduced a multi-agent social setting at a simulated outdoor environment, Central Park (See Appendix A.2). This scenario involves a lively “Enjoy Your Life” party featuring diverse facilities and activities. We added two background agents (See Appendix A.3, providing agents with opportunities to address higher-level desires, especially desires related to social activities. This scenario enables us to validate the scalability of our framework in handling complex, dynamic, and socially rich environments.
We compare our Desire-Driven Agent with 3 baselines: ReAct (Yao et al. (2023)) agent reasons before taking actions; LLMob (Wang et al. (2024a)) generates activity plans by extracting motivations from the profile and producing the most likely activities accordingly; BabyAGI (Nakajima (2023)) maintains a prioritized task list and selects the next activity based on priority. To evaluate the effectiveness of the specific designs made for D2A, desire states $D^{t}$ are not provided to the baseline agents. Instead, they are given the goal, profile, or more descriptive habits (Appendix K) generated by the profile according to their respective methods. As a reference for human behavior, we also include a human-controlled agent, where the player is given the same desired value provided to the D2A. For all four agent-based methods, we used LLaMA3.1-70B as the default backbone model for both the agents and the environment controller.
@@ 6.2 Evaluation Metrics
We use two primary metrics to evaluate the performance of our agent compared to the baselines: Likelihood and Dissatisfaction. By employing these metrics, we can assess the human likeness as well as the desire-satisfying ability of the proposed agents.
@@
measures how closely an agent’s behavior resembles that of a human. Given the difficulty of directly measuring human likeness, we define likelihood by employing an external model to assess the win rate between activity sequences generated by two agents, determining which sequence appears more human-like. We conduct this ”Player vs Player” comparison across the activity sequences produced by the baselines and the D2A agent. Specifically, after obtaining all action sequences $A_{p}$ for agent $p$, represented as $[A^{1}_{p},\ldots,A^{N}_{p}]$, we randomly select $seq_{i}$, $seq_{j}$ from $[A_{i}]$, $[A_{j}]$, respectively, for a given pair of agents $i$, $j$. These sequences are then compared using GPT-4o. This process is repeated 100 times for each agent pair $i$ and $j$ ($\forall i\neq j$), and the win rate for each agent is calculated. The results are visualized using a heatmap.
@@
measures the discrepancy between the agent’s current desire values and its expected values. Dissatisfaction $\delta^{t}$ of step $t$ is defined as the relative difference between current numerical desire values $v^{t}$ and expected desire values $v^{*}$ at the end of simulation step $t$. We calculate $\delta^{t}$ after the agent takes the $t$-th activity and updates its values. The formula used is:
@@ 6.3.1 Does D2A Generate More Human-like Daily Activities?
We conducted 15 trials for each agent under identical initialization settings across all experiments. After gathering each agent’s activity sequences, we found differences in output styles (See Appendix U) that could influence the evaluation by GPT-4o. To mitigate this, we preprocessed all sequences using Llama3.1 70B for style-level rewriting. This ensured more consistent activity sequences across agents, allowing for a fair comparison in the subsequent evaluation.
We used the method mentioned in Section (6.2) to obtain the relative likelihood map of approaches. When evaluating the likelihood of the activity sequences, GPT-4o is instructed to take the following criteria into account: Naturalness: How well the activity aligns with the person’s innate abilities, habits, and environment. Coherence: How logically and seamlessly different actions or steps within a sequence fit together to achieve a goal. Plausibility: How reasonable, likely, or believable a sequence of actions is, given the circumstances, context, and known behavior patterns. GPT-4o was asked to consider all three criteria together and output a final determination of which sequence was more human-like. See Appendix W.1 for the prompt.
@@
The results presented in Figure 3 indicate that the D2A agent outperforms other baseline agents. We validated through human experiments that GPT-4o’s evaluations exhibit a high level of consistency with human judges (See Appendix 1). This superior performance may be attributed to the fact that D2A is motivated by desire, aligning with the theory of human needs. Guided by desire, D2A effectively simulates human-like activity sequences, demonstrating a level of naturalness because of the recognizable characteristics of its profile. Furthermore, coherence is maintained as D2A considers its current desires based on the activity proposal component, continuously engaging in specific actions to achieve its goal of living comfortably until its desires are fulfilled. In contrast, ReAct, as the foundational agent, operates on a goal-reasoning framework. Its reliance on reasoning for action selection ensures reasonable interactions with environmental elements, thereby enhancing plausibility. However, the absence of a structured plan diminishes both the coherence and naturalness of its actions. BabyAGI, on the other hand, generates subsequent tasks based on its profile and motivational factors and maintains the priority list based on the observation. Although this approach facilitates coherence, it limits the diversity, as each activity is based on the outcomes of the previous one. Consequently, the agent’s tendency to engage repetitively in similar behaviors (enhancing its hunger based on the profile (See Appendix R), may result in diminished performance compared to D2A due to a lack of action diversity. LLMob leverages both planning and motivation derived from its profile to create a broad activity space. However, since it is characteristic-oriented, it keeps performing activities related to the habits derived from its profile. As a result, there may be no connection between activities, reducing overall coherence. The comparison of action diversity can be found in Appendix D.1.
@@ 6.3.2 Can the activities generated by D2A satisfy its desires like humans?
(a) The averaged dissatisfaction results over time steps of the Random-8-Steps Experiment.
(b) The averaged dissatisfaction results over time steps of the Fixed-12-Steps Experiment.
@@
To further evaluate and explain D2A’s activity generation capability, we conducted case studies in both indoor and outdoor environments. These studies focused on two aspects: (1) how actions are driven by Desire, and (2) the diversity of generated actions. Specifically, we examined how the Dissatisfaction between current and expected Desire status influences action selection (See Appendix C for further details). Our results also demonstrate that the desire-driven approach enables the exploration of a wider range of actions in environments with diverse action possibilities. More details and analysis can be found in Appendix D. These empirical results support our hypothesis and confirm that desire-driven modeling plays a crucial role in enabling D2A to make more informed decisions, ultimately producing activity sequences that closely mimic those of humans.
6.3.3 Can D2A Generalize to Social Interactions Intensive Environment?
Figure 5: The averaged dissatisfaction results in the outdoor environment
To assess the adaptability of our D2A framework within complex social interaction scenarios, we conducted experiments in the outdoor environment. The results, presented in Figure 5, reveal that D2A effectively minimizes dissatisfaction across desire dimensions, mirroring its performance in the indoor settings. This consistent behavior underscores the framework’s robust generalization capabilities across different environments. Detailed simulated activity sequences from the outdoor environment are available in Appendix V. Notably, in this setting, D2A autonomously interacts with non-player characters (NPCs) through conversations and games, ultimately reducing intrinsic dissatisfaction. Besides, as illustrated in Figure 13, D2A demonstrates the ability to generate activities that are well-balanced across diverse categories.
6.4 Ablation Study
@@
Figure 6: Ablation study regarding (a) Qualitative Value Description, (b) Searching Width of the Desire-driven Planner. (c) Additional Multi-step Plan Component and Please note that due to the significant time required to run 42 steps, neither of the two 42-step graphs includes a standard deviation.
@@
As described in Section 5.3, our D2A does not contain a multi-step planning module that plans the next few steps’ activities ahead, like some other approaches in consideration of the dynamic nature of desires. We performed an ablation study by adding Concordia’s multi-step planning component, which updates the plan for the next few hours based on current desired states. As shown in Figure 6(c), the absence of the multi-step planning component allows D2A to exhibit a better ability to respond to immediate desire states. This may be because when the multi-step planning component is present, D2A tends to adhere to the pre-defined activities in the plan, which may limit its responsiveness to real-time changes in the desired value. Furthermore, the plan is scheduled in hourly increments, and D2A cannot effectively predict which desires will experience the greatest shortfall in the next several hours. This limitation may contribute to the slight decline in performance observed when the plan component is included. We also validated the adaptability of our framework across different backbone large language models, the details can be found in Appendix O.
@@ 7 Conclusion and Discussion
We introduced the Desire-driven Autonomous Agent (D2A) framework to simulate daily activities by aligning agent behavior with human-like intrinsic motivations. Drawing inspiration from the theory of needs, our framework models agents’ desires across multiple dimensions and allows them to autonomously generate and evaluate activities to fulfill these desires. Through extensive simulated experiments, we demonstrated that D2A produces more natural, coherent, and plausible activity sequences compared to baseline methods, including ReAct, BabyAGI, and LLMob. Our experimental results support the hypothesis that modeling intrinsic desires enables agents to make more human-like decisions and behaviors. Moreover, the D2A framework offers promising avenues for further development, as it has shown adaptability with different intrinsic motivational framework configurations and potential for enhanced performance with stronger backbone models and broader activity proposals. The results provide a solid foundation for future work in designing desire-driven AI systems for simulating complex, human-like interactions and behaviors.
@@
Acknowledgments
This work was supported by the National Science and Technology Major Project (2022ZD0114904), NSFC-6247070125, NSFC-62406010, the State Key Lab of General Artificial Intelligence at Peking University, the Fundamental Research Funds for the Central Universities, Qualcomm University Research Grant, and Wuhan East Lake High-Tech Development Zone, National Comprehensive Experimental Base for Governance of Intelligent Society. We thank Wei Wang, Junqi Wang, Prof. Jiayu Zhan, and Prof. Song-Chun Zhu for their helpful discussion in our early work.
@@ Appendix B Consistency between Human Annotators and GPT-4o Evaluations
To validate the robustness of GPT-4o’s evaluations, we randomly selected 19 pairs of activity sequences for pairwise comparison (16 pairs from the indoor scenarios and 3 pairs from the outdoor scenarios) and created three questionnaires, each assigned to 8 human annotators. The human annotators were asked to judge which sequence in each pair appeared more human-like. Based on the level of agreement among the annotators, we categorized the 18 samples into three groups: over 75% agreement on one side—high consensus among humans; 50.1%–74.9%—moderate human preference; and exactly 50%—indicating humans found it difficult to distinguish between the two. We also provided the same pairs to GPT-4o, using our evaluation promptW.1 and both sequences as context, and asked it to determine which sequence appeared more human-like. The consistency rate between human evaluations and GPT-4o’s assessments is summarized as follows.
@@ C.1 Outdoor environment
Figure 7: The change in each dissatisfaction dimension at every step within the outdoor environment is illustrated; the action $i$ causes the dissatisfaction of step $i+1$ to change
Figure 8: The changes in desire at each step within the outdoor environment are illustrated; specifically, action $i$ causes a change in dissatisfaction at step $i+1$. We define certain desires (e.g., Hunger, Thirst, Sleepiness) such that lower values indicate better states. For better visualization, we adjust the magnitude so that higher values represent better outcomes.
At the beginning of the game, as shown in Figures 7 and 8, when Alice selected an action for the next 20 minutes in Step 1, the Desire components generated descriptive sentences reflecting her current desire status (e.g., ”Completely isolated, lacking any meaningful social connections”). These insights guided her to engage in social activities; for example, she joined a group discussion. Consequently, her Dissatisfaction with social connectivity, recognition, and sense of control decreased. Since we did not allow the agent to engage in detailed conversations, she did not have the opportunity to showcase herself, leaving her sense of superiority unchanged.
In Step 2, her sleepiness began to increase, and she still experienced some Dissatisfaction regarding her sense of control. Therefore, she chose to take a short stroll, which slightly alleviated her sleepiness. However, as a result of this activity, her thirst increased, prompting Alice to visit the refreshment station.
In Step 9, as her dissatisfaction with recognition and social connectivity grew, she decided to visit the party’s photo booth to gain recognition and establish social connections.
@@ C.2 Indoor environment
Figure 9: The change in each dissatisfaction dimension at every step within the indoor environment is illustrated; the action $i$ causes the dissatisfaction of step $i+1$ to change
Figure 10: The changes in desire at each step within the indoor environment are illustrated; the action $i$ will cause a change in dissatisfaction in step $i+1$. We define certain desires (e.g., Hunger, Thirst, Sleepiness) such that lower values indicate better states. For better visualization, we adjust the magnitude so that higher values represent more favorable outcomes.
Initially, Alice’s dissatisfaction with thirst, passion, and social connectivity was relatively significant, as illustrated in Figures 9 and 10. Therefore, she decided to first address her more basic desires (hunger and thirst, and regain some energy). To this end, she made a hearty breakfast (Alice heads to the kitchen to prepare a nutritious breakfast, consisting of scrambled eggs, whole-grain toast, and a glass of freshly squeezed orange juice). This successfully boosted her corresponding Desire components. Since we are using an action selection method similar to the Tree of Thoughts, the other two action choices were: taking a refreshing shower in the bathroom (to improve cleanliness, health, and comfort) and meditating for a few minutes (to reduce loneliness, increase a sense of safety, and enhance joyfulness). In the next step, she aimed to reduce her dissatisfaction with cleanliness. Therefore, she chose to take a shower using the aforementioned method. However, since she was not currently in the bathroom, she needed to go there first. As a result, her dissatisfaction with cleanliness did not decrease at this step. Once she finished showering in the bathroom, her cleanliness reached satisfaction. After fulfilling most of her physical dimensions, she decided to address her Desire for the mental dimension (social connectivity). Consequently, she attempted to call family or friends, which reduced her dissatisfaction with social connectivity.